OpenStudy (sushi121212):

Factor: 200=(30-2x)(50-2x)

OpenStudy (vuriffy):

Do you know how to multiply terms in factoring?

OpenStudy (sushi121212):

Yes

OpenStudy (sushi121212):

So it's 200=1500-160x-4x^2?

OpenStudy (sushi121212):

What do I do after that though?

OpenStudy (vuriffy):

200=4x^2 - 160x + 1500.

OpenStudy (vuriffy):

You may subtract 1500. Which results to be: ?

OpenStudy (sushi121212):

1300=4x^2 - 160x

OpenStudy (vuriffy):

I believe you would need everything to one side. So you can just move that 1300 over. 4x^2 - 160x - 1300. Do you need the x values, yes?

OpenStudy (sushi121212):

So do I simplify it to x^2-40-325?

OpenStudy (sushi121212):

Oops, I meant x^2-40x-325

OpenStudy (vuriffy):

You may, that will not change the x values. Are you trying to find x?

OpenStudy (vuriffy):

But, if your just simplifying, that is the lowest it can go.

OpenStudy (sushi121212):

Yes

OpenStudy (sushi121212):

I'm trying to find x.

OpenStudy (vuriffy):

Are you familiar with the quadratic formula?

OpenStudy (vuriffy):

\[-b +/- \sqrt{b^2 - 4ac} /(2a)\]

OpenStudy (sushi121212):

I don't think I've learned it yet.

OpenStudy (vuriffy):

One second, sorry.

OpenStudy (sushi121212):

It's okay, take your time.

OpenStudy (vuriffy):

a = 1 b = -40 c = -325. x^2-40x-325 Do you understand why it is these terms. (1)x^2 (-40)x (-325)

OpenStudy (sushi121212):

Yes

OpenStudy (sushi121212):

I'm so sorry, but I have to go now... It's okay if you don't finish solving it! I appreciate your help!!

OpenStudy (vuriffy):

I apologize I kept getting errors on my calculator, not sure what I messed up on, because I figured out the x value a different method, but it's more complex than this.

OpenStudy (vuriffy):

But look at this,

OpenStudy (vuriffy):

\[x = -5 (\sqrt{29} -4) \]

OpenStudy (vuriffy):

\[x = 5(4+\sqrt{29})\]

OpenStudy (vuriffy):

You may distribute the 5 and -5 to both values and get the exact decimal.

OpenStudy (sushi121212):

Okay, thank you so much!!!